### v2

@@ 4 Discussion and Conclusion
We assess the fidelity of SFNO-BVMC on overall ensemble diagnostics, spectral diagnostics, and extremes diagnostics. This comprehensive pipeline is specifically designed for ensemble forecasts (not solely for deterministic ones). As the field of ML-based ensemble forecasting rapidly grows, we hope that other groups also use these statistics to evaluate their ensembles. On overall diagnostics, SFNO-BVMC’s performance is 18 hours behind IFS ENS, a leading operational weather forecast on most diagnostics and most variables. We present a pipeline to evaluate the ensemble’s performance on extreme 2m temperature events.
@@
The spectral diagnostics demonstrate that individual ensemble members in the SFNO-BVMC have blurry predictions compared to ERA5. We minimize this as much as possible through a small scale factor, a large embedding dimension, and no autoregressive fine-tuning. Still, some degree of blurring still remains. However, our spectral diagnostics reveal that the spectra from SFNO-BVMC remain constant throughout the rollout. This means that SFNO-BVMC’s ability to predict extreme weather and fine-scale phenomena remains constant. Additionally, the SFNO-BVMC ensemble-mean spectra indicate that the ensemble members realistically diverge. Future research and architectural improvements are necessary to reduce the extent of the initial blurring.
@@
Bred vectors are open-sourced through the earth2mip package, and they can readily be applied to other deterministic architectures. This enables out-of-the-box ensemble forecasts from the wide array of existing deterministic architectures. Indeed, recently, there have been over twenty deterministic ML weather prediction models (Arcomano et al., 2020; Bi et al., 2023; Nguyen et al., 2023; Chen et al., 2023b; Bodnar et al., 2024; Mitra and Ramavajjala, 2023; Ramavajjala, 2024; Pathak et al., 2022; Bonev et al., 2023; Weyn et al., 2021; Willard et al., 2024; Keisler, 2022; Karlbauer et al., 2023; Rasp et al., 2024, 2020; Lang et al., 2024; Couairon et al., 2024; Scher and Messori, 2021; Chen et al., 2023a). It is computationally expensive and programmer time-intensive to convert all these architectures into ensembles using probabilistic training (e.g. through diffusion models or through CRPS training). Even for the architectures that are converted to probabilistic training, bred vectors and multiple checkpoints can provide baseline ensemble scores. This baseline can be used to guide further development of end-to-end training.
@@
Understanding how ML models respond to perturbations is an important research frontier (Bülte et al., 2024; Selz and Craig, 2023). Some initial condition perturbations rely on external sources. For instance, Price et al. (2023) use perturbations from operational data assimilation that provide valuable information about observational uncertainty. With the PanGu ML model, Bülte et al. (2024) create ensembles using IFS perturbations, but they find that these perturbations do not lead to the best performance. Other initial condition perturbations, such as bred vectors, do not rely on external sources. If an ML model is used to emulate climate models (e.g. in Watt-Meyer et al. (2023)), these perturbations are still available, unlike IFS or data assimilation perturbations. Additionally, there are often a limited number of external perturbations from existing weather center. However, bred vectors can be used to generate arbitrarily large ensembles, such as the huge ensemble in Part II. For operational weather forecasting, future work is necessary to improve calibration by combining multiple types of perturbations.
@@
Looking to the future of ML-based ensemble forecasting, an important design choice is whether the ensemble is created during training or after training. NeuralGCM (Kochkov et al., 2023) and GenCast (Price et al., 2023) create ensembles end-to-end during training; they train using probabilistic loss functions. Here, we train SFNO using a deterministic loss function, and we create the ensemble after training. In the machine learning literature, it is an open area of research whether ensemble training or post hoc ensembling leads to the most reliable results (Jeffares et al., 2023). In weather forecasting, so far, GenCast and NeuralGCM offer superior ensemble performance to SFNO-BVMC. They have better CRPS scores and spread-skill ratios. Even at full ERA5 horizontal resolution, GenCast does not produce blurry forecasts. While GenCast and SFNO-BVMC run on different hardware (TPUs, compared to NVIDIA GPUs used here), GenCast takes 6 minutes to create a 2-week forecast, with a timestep of 12 hours. At the same horizontal resolution, SFNO-BVMC takes 1 minute to create a 2-week forecast, with a timestep of 6 hours; therefore, SFNO-BVMC appears to be a factor of 12 faster for inference. In part, this difference is because SFNO-BVMC does not require the iterative denoising used by GenCast at each timestep. In Part II of this paper, we assess the performance of huge ensembles of SFNO-BVMC. A promising area of future research is to explore the behavior of huge ensembles from these other ML-based models.
@@
The current generation of ML-based ensemble weather forecasts all have core design differences. IFS ENS uses physics-based modeling, NeuralGCM uses a differentiable dynamical core and an ML physics parameterization, GenCast uses a diffusion-based generative model, and SFNO-BVMC uses deterministic training. Because of these differences, future research is necessary to assess the strengths and weaknesses of each model in different meteorological regimes. When different forecasting systems have uncorrelated errors, a multimodel ensemble can lead to improved skill. Each forecasting system could be post-processed, bias-corrected, and optimized to create the best ensembles for each region.
@@
In this manuscript, we run our extreme diagnostics pipeline on warm temperature extremes, and we validate on summer 2023, as it is the hottest summer in the observed record. However, future work is necessary to characterize the performance on cold temperature extremes and other extreme events of interest (e.g. extreme wind, vapor transport, or precipitation). The EFI here is calculated on daily mean temperature, but it can also be calculated for other quantities, such as daily max or min temperature, convective available potential energy, or vapor transport (Lavers et al., 2016). Similarly, the ROC curves and reliability diagrams could be calculated for other types of extremes. We have presented an ensemble extreme diagnostics pipeline that can be used to guide development for other ML data-driven weather systems.
In Part II, we use SFNO-BVMC to generate a huge ensemble, with 7,424 members. This ensemble is 150x larger than the ensembles used for operational weather forecasting. We explore how an ensemble of this size enables analysis of the statistics of low-likelihood, high-impact extremes.
Appendix A Perturbation Amplitudes
@@
Appendix B Definition of Spread and Error
@@
Appendix C Reliability Diagrams for Additional Lead Times
@@ Acknowledgements.
The code, datasets, and models are all stored at https://doi.org/10.5061/dryad.2rbnzs80n. The code is integrated with Zenodo: https://doi.org/10.5281/zenodo.14710345. We include the code to train SFNO, conduct ensemble inference with bred vectors and multiple checkpoints, and scoring and analysis code. We also open-source the model weights of the trained SFNO. See the README of the former DOI for information on how to use the codebase and for the permissive license associated with the code and data.

### v3

@@ 4 Discussion and Conclusion
We assess the fidelity of SFNO-BVMC on overall ensemble diagnostics, spectral diagnostics, and extremes diagnostics. This comprehensive pipeline is specifically designed for ensemble forecasts (not solely for deterministic ones). As the field of ML-based ensemble forecasting rapidly grows, we hope that other groups also use these statistics to evaluate their ensembles. On overall diagnostics, SFNO-BVMC’s performance lags approximately 12–18 hours behind IFS ENS. We present a pipeline to evaluate the ensemble’s performance on extreme 2m temperature, 10m wind speed, and heat index events.
@@
The spectral diagnostics demonstrate that individual ensemble members in the SFNO-BVMC have blurry predictions compared to ERA5. We minimize this as much as possible through a small scale factor, a large embedding dimension, and no autoregressive fine-tuning. Still, some degree of blurring still remains. However, our spectral diagnostics reveal that the spectra from SFNO-BVMC remain constant throughout the rollout. Additionally, the SFNO-BVMC ensemble-mean spectra indicate that the ensemble members realistically diverge. Future research and architectural improvements are necessary to reduce the extent of the initial blurring.
@@
Bred vectors are open-sourced through the earth2mip package, and they can readily be applied to other deterministic architectures. This enables out-of-the-box ensemble forecasts from the wide array of existing deterministic architectures. Indeed, recently, there have been over twenty deterministic ML weather prediction models (Arcomano et al., 2020; Bi et al., 2023; Nguyen et al., 2023; Chen et al., 2023b; Bodnar et al., 2024; Mitra and Ramavajjala, 2023; Ramavajjala, 2024; Pathak et al., 2022; Bonev et al., 2023; Weyn et al., 2021; Willard et al., 2024; Keisler, 2022; Karlbauer et al., 2023; Rasp et al., 2024, 2020; Lang et al., 2024; Couairon et al., 2024; Scher and Messori, 2021; Chen et al., 2023a). It is computationally expensive and programmer time-intensive to convert all these architectures into ensembles using probabilistic training (e.g. through diffusion models or through training on the CRPS loss function). Even for the architectures that are converted to probabilistic training, bred vectors and multiple checkpoints can provide baseline ensemble scores. This baseline can be used to guide further development of end-to-end training.
@@
Understanding how ML models respond to perturbations is an important research frontier (Bülte et al., 2024; Selz and Craig, 2023).Understanding how ML models respond to perturbations is an important research frontier (Bülte et al., 2024; Selz and Craig, 2023). In particular, future work is necessary to compare the computational cost and skill of different initial condition perturbation methods (Bülte et al., 2024), in tandem with model perturbations. We find that bred vectors are a computationally inexpensive way to achieve reasonable spread-error ratios and to generate an arbitrarily large ensemble. Further refinement of initial condition perturbation techniques is needed to improve forecast performance. Two advantages of bred vectors are that they do not rely on external sources, and they can be used to generate arbitrarily large ensembles. First, Price et al. (2023) use external perturbations from operational data assimilation to include estimates of observational uncertainty. With the PanGu ML model, Bülte et al. (2024) test ML ensembles with IFS perturbations. Bred vectors, do not rely on external sources. If an ML model is used to emulate climate models (e.g. in Watt-Meyer et al. (2023)), bred vector perturbations are still available, unlike IFS or data assimilation perturbations. Second, there are often a limited number of external perturbations from existing weather centers. Bred vectors can be used to generate arbitrarily large ensembles, such as the huge ensemble in Part II.
@@
Looking to the future of ML-based ensemble forecasting, an important design choice is whether the ensemble is created during training or after training. NeuralGCM (Kochkov et al., 2023) and GenCast (Price et al., 2023) create ensembles end-to-end during training; they train using probabilistic loss functions. Here, we train SFNO using a deterministic loss function, and we create the ensemble after training. In the machine learning literature, it is an active area of research whether ensemble training or post hoc ensembling leads to the most reliable results (Jeffares et al., 2023). In weather forecasting, so far, GenCast and NeuralGCM offer superior ensemble performance to SFNO-BVMC. They have better CRPS scores and spread-skill ratios. Even at full ERA5 horizontal resolution, GenCast does not produce blurry forecasts. While GenCast and SFNO-BVMC run on different hardware (TPUs, compared to NVIDIA GPUs used here), GenCast takes 6 minutes to create a 2-week forecast, with a timestep of 12 hours. At the same horizontal resolution, SFNO-BVMC takes 1 minute to create a 2-week forecast, with a timestep of 6 hours; therefore, SFNO-BVMC appears to be a factor of 12 faster for inference. In part, this difference is because SFNO-BVMC does not require the iterative denoising used by GenCast at each timestep. In Part II of this paper, we assess the performance of huge ensembles of SFNO-BVMC. A promising area of future research is to explore the behavior of huge ensembles from these other ML-based models.
@@
The current generation of ML-based ensemble weather forecasts all have core design differences. IFS ENS uses physics-based modeling, NeuralGCM uses a differentiable dynamical core and an ML physics parameterization, GenCast uses a diffusion-based generative model, and SFNO-BVMC uses deterministic training. Because of these differences, future work is necessary to assess the strengths and weaknesses of each model in different meteorological regimes. When different forecasting systems have uncorrelated errors, a multimodel ensemble can lead to improved skill. Each forecasting system could be post-processed, bias-corrected, and optimized to create the best ensembles for each region.
@@
In this manuscript, we run our extreme diagnostics pipeline on warm temperature extremes, and we validate on summer 2023, as it is the hottest summer in the observed record. At lead times of 48 hours and 96 hours, the performance on cold extremes and wind extremes is similar to IFS. However, future work is necessary to reduce false positives for these other classes of extremes at 10-day lead times (Figure 20). The EFI here is calculated on daily mean temperature, but it can also be calculated for other quantities, such as daily max or min temperature, convective available potential energy, or vapor transport (Lavers et al., 2016). Similarly, the ROC curves and reliability diagrams could be calculated for other types of extremes. We have presented an ensemble extreme diagnostics pipeline that can be used to guide development for other ML data-driven weather systems.
In Part II, we use SFNO-BVMC to generate a huge ensemble, with 7,424 members. This ensemble is 150x larger than the ensembles used for operational weather forecasting. We explore how an ensemble of this size enables analysis of low-likelihood, high-impact extremes.
Appendix A Case Study: 2023 Phoenix Heatwave
We include a case study for a heatwave in Phoenix in summer 2023. Phoenix had temperatures above 310 K (36.85 C) for over 30 consecutive says in summer 2023. We compare the ensemble forecasts from SFNO-BVMC and the IFS ensemble in Figure 16. We compare SFNO-BVMC to the IFS ensemble, and we visualize their respective verification datasets. We show that at a lead time of 3 days, SFNO-BVMC can forecast the high temperatures observed over the region during the region. The IFS ensemble is initialized with an operational analysis, not ERA5, and we use this analysis as the verification dataset for IFS. Notably, the operational analysis has even sharper fields than ERA5: this has previously been quantified in Figure S38 and Supplementary Materials Section 7.5.3 of Lam et al. (2023). The difference between operational analysis and ERA5 reanalysis is shown for this heatwave in Figure 16.
Figure 16: 2023 Phoenix Heatwave. Comparison of ensemble forecasts from SFNO-BVMC (a) and the IFS ensemble (b) at a grid cell near Phoenix, Arizona, USA. Both models’ forecasts are initialized on June 27, 2023 at 00:00 UTC. The SFNO-BVMC verification dataset is ERA5 and the IFS ENS verification dataset is operational analysis. (c) and (e) show ERA5 and operational analysis, respectively, for the daily max temperature on June 30, 2023. (d) and (f) show the SFNO-BVMC ensemble mean and IFS ENS mean, respectively, for the daily max temperature on June 30, 2023. The black stars in c-f denote the grid cell near Phoenix, Arizona, USA used in (a) and (b).
Appendix B Perturbation Amplitudes
@@
Appendix C Definition of Spread and Error
@@
Appendix D Diagnostics for Additional Variables and Lead Times
@@
We also include the reliability diagrams for the heat index (Lu and Romps, 2022), which combines 2m temperature and moisture, and for 10m wind speed. In Figure 19, we show the reliability diagrams for 95th percentile heat index. As for 2m air temperature, the 95th percentile heat index is calculated from 1993-2016 ERA5 climatology. Comparing Figure 19 to Figure 15, we find that SFNO-BVMC is similarly reliable for heat index extremes as it is for warm 2m temperature extremes.
Figure 19: Reliability Diagram for Heat Index. The heat index combines 2m air temperature and 2m dewpoint. Reliability diagrams are shown for 95th percentile heat index extremes at a lead time of 48 hours and 98 hours (a) and 120 and 240 hours (b). Reliability diagrams are calculated using summer 2023 forecasts, from June 1, 2023 to Aug 14, 2023.
Next, we show overall CRPS scores and reliability diagrams for 10m wind speed, and we also show the model’s reliability for forecasting cold extremes (5th percentile 2m temperature). To calculate these statistics, we use December-January-February (DJF) from 2021 to 2022, while we use summer 2023 in Figure 15. We use a winter season for the Northern Hemisphere to include sufficiently cold extremes over land in the Northern Hemisphere. Additionally, DJF 21-22 is within the time period included by the WeatherBench dataset (Rasp et al., 2024), so we can readily access the IFS ensemble wind forecasts via a Zarr file stored on Google Cloud, without having to download additional data from ECMWF’s tape servers. On 10m wind speed and cold extreme air temperature, SFNO-BVMC and the IFS ensemble have similar performance on overall CRPS and on reliability diagrams (Figure 20). This indicates that the ensemble generation methodology (bred vectors and multiple checkpoints) is promising for other variables and classes of extremes. SFNO-BVMC does degrade in reliability when making forecasts of extreme wind with 90% probability. This problem is accentuated at longer lead times (see below), and future research is necessary to isolate the cause of this behavior.
Figure 20: SFNO-BVMC Performance on 10m Wind Speed and Cold Extremes. (a) Overall CRPS for SFNO-BVMC and the IFS ensemble on 10m wind speed. Lower CRPS scores are better (b) Reliability Diagrams for 95th Percentile 10m wind speed for SFNO-BVMC and IFS at lead times of 48 and 96 hours. (c) Reliability Diagrams for 5th percentile temperature extremes for SFNO-BVMC and IFS at 48 and 96 hour lead times of 48 and 96 hours. All scores are calculated using all forecasts initialized in December-January-February 2021. Successful forecasts lie along the 1-to-1 line.
We identify two areas for future research and model improvement. First, interestingly, both SFNO-BVMC and the IFS ensemble have degraded performance for forecasting cold extremes, as opposed to warm extremes (compare the 5th percentile reliability in Figure 20 to the 95th percentile reliability diagrams in Figure 18). With future model development, we hope to improve the performance of SFNO-BVMC in forecasting cold extremes. Second, for 10m wind speed and cold temperature extremes at a lead time of 10 days, SFNO-BVMC’s reliability degrades (Figure 21). For these variables, the ensemble still has a good overall forecast scores (see the wind speed CRPS in Figure 20 and the 2m temperature Figure 7). SFNO-BVMC reliability is close to IFS for extreme forecast probabilities from 0% to 50%. However, the reliability drops when the model predicts high probabilities (greater than 70%) of extreme conditions. In these cases, SFNO-BVMC tends to be overconfident: its forecast of an extreme event does not match the observed outcome. This overconfidence occurs extremely rarely. At a lead time of 10 days, it is very uncommon (less than 1% of all forecasts for 2m temperature, less than 0.1% for 10m wind speed) for SFNO-BVMC to predict a greater than 70% chance of extreme wind or cold temperatures. At this long lead time, there is significant ensemble spread induced by the perturbations, so the ensemble system is not confident in issuing extreme forecasts. Still, having a calibrated reliability diagram is crucial for all forecast probabilities, and this shortcoming must be resolved with future model development.
Figure 21: Reliability Diagram at 240h Lead Time for 95th percentile wind speed and 5th percentile temperature extremes. Reliability diagrams are shown for 95th percentile extremes at a lead time of 48 hours and 98 hours. Reliability diagrams are calculated using all forecasts initialized in December-January-February 2021. Successful forecasts lie along the 1-to-1 line.
Figure 22: Perturbed Spectra at 0h. Same as Figure 11, but showing the spectra of perturbations applied to the ERA5 initial conditions.
Appendix E Solar Zenith Error in Computer Code When Calculating Initial Condition Perturbations
After performing the analysis in this manuscript, we discovered an error in our calculation of the initial condition perturbations. During the first steps of calculating bred vectors at $t_{-2}$ and $t_{-1}$ (Figure 4), we incorrectly supplied SFNO with the solar zenith angle at time $t_{0}$. We have verified and established that this error does not affect any of the conclusions or scores presented in this manuscript. This error does not make a discernible difference for three reasons. First, the last breeding step calculates the perturbations using SFNO at $t_{0}$. At $t_{0}$, the correct zenith angle is supplied, so the final perturbation is still based on the correct SFNO forecasts. Second, we validate that the error does not cause undesired artifacts related to the diurnal cycle in the actual perturbations (see Figure 5). Third, the breeding cycle only uses 1-step forecasts, which means that the error from using the incorrect zenith angle does not grow.
We also note that this error does not affect the 15-day rollout of SFNO, only the calculation of the bred vectors. Given the nature of this error, we do not believe it would cause SFNO-BVMC to appear better than it actually is. Instead, it would be more likely to degrade its performance, making the method seem worse than it really is. We compare the ensemble scores for the ensemble with the error (named "Original") and the fixed ensemble (named "Fixed") in Figures 23 and 24; the scores are nearly identical. We have corrected the error in the GitHub page for our project, but for scientific reproducibility, the error remains in the codebase in the DOI in our Code and Data Availability Section, since this is the version of the code that we used for our analysis.
Figure 23: CRPS Comparison of Original and Fixed Bred Vector Perturbation Method. Our original calculation of bred vectors contained an error with a mismatched cosine zenith angle during the first two breeding steps. This figure compares the CRPS of an ensemble with the "Original" (incorrect) bred vector calculation and the "Fixed" calculation for 2m and 850hPa atmospheric temperatures, the 500hPa geopotential height, and the 850hPa zonal wind as representative fields. Results are shown for 52 forecasts initialized in summer 2020, one per week starting January 2.
Figure 24: Ensemble Mean RMSE and Ensemble Spread Comparison of Original and Fixed Bred Vector Perturbation Method. Our original calculation of bred vectors contained an error with a mismatched cosine zenith angle during the first two breeding steps. This figure compares the ensemble mean RMSE and ensemble spread of an ensemble with the "Original" (incorrect) bred vector calculation, and the "Fixed" calculation. Results are shown for 52 forecasts initialized in summer 2020, one per week starting January 2.
@@ Acknowledgements.
The code, datasets, and models are all stored at https://doi.org/10.5061/dryad.2rbnzs80n, via DataDryad. The code is integrated with Zenodo at the prior DOI. We include the code to train SFNO, conduct ensemble inference with bred vectors and multiple checkpoints, and scoring and analysis code. We also open-source the model weights of the trained SFNO. See the README of the DOI for information on how to use the codebase and for the permissive license associated with the code and data. The code is available via the Lawrence Berkeley Lab BSD variant license, and the data is available with a CC0 license. For convenience, the webpage of our project is https://github.com/ankurmahesh/earth2mip-fork.